OpenStudy (anonymous):

On January 1, 2010, Chessville has a population of 50,000 people. Chessvill then enters a period of population growth. Its population increases 7% each year. On the same day, Checkersville has a population of 70,000 people. Checkersville starts to experience a population of decline. Its population decreases 4% each year. During what year will the population of chessville first exceed that of Checkersville? Will you explane how you did it as well? I give medals and fandom...

OpenStudy (anonymous):

@frogs29 Help?

OpenStudy (anonymous):

@dirtrider do you know?

OpenStudy (anonymous):

hold on ill figure it out

OpenStudy (anonymous):

figure it out yet?

OpenStudy (anonymous):

50000*(1+0.07)^y=70000*(1-0.04)^y --> y=3.1 (years) that's when they'll be equal so during the third year Chessville will have a greater population

OpenStudy (anonymous):

sorry i got busy

OpenStudy (anonymous):

fyi : these type of problems are exponential growth/decay problems so look up the formula's for those

OpenStudy (anonymous):

@frogs29 its ok :)

OpenStudy (anonymous):

xjgnui dvjkxsmgn